# Energetic particles accelerated via turbulent magnetic reconnection in protoplanetary discs – I: ionisation rates

### v1

@@ Abstract
Context. Ionisation controls the chemistry, thermal balance, and magnetic coupling in protoplanetary discs. However, standard ionisation vectors such as stellar UV, X-rays, Galactic Cosmic Rays might not be efficient enough, as UV/X-rays are attenuated rapidly with depth, while Galactic Cosmic Rays are modulated. Turbulence-induced magnetic reconnection in disc atmospheric layers offers a physically motivated, in-situ source of energetic particles (EPs) that has never been considered.
@@
Methods. We adopt a Fermi-like acceleration model with parameters linked to a turbulent reconnection geometry trigger by the magneto-rotational instability, yielding a steady-state energy distribution of the EP forming a power-law of index $p=2.5$. We propagate electrons and protons through the disc and compute primary and secondary ionisation and associated heating on a fiducial T Tauri disc model background. The non-thermal normalisation is set by the fraction of local viscous accretion energy dissipation channelled to EPs, parametrized by $\kappa$.
Results. For $\kappa\gtrsim 0.4\%$, EPs ionisation overpass standard sources like X-rays and galactic cosmic rays in the disc atmosphere and intermediate/deep layers out to radii of a few tens of AU. Even at $\kappa\sim 0.025\%$, EPs contribute at the few-percent level, thus are chemically and dynamically relevant. EP-induced heating complements UV/X-ray heating in the atmosphere and persists deeper. These results identify EPs accelerated by turbulence-induced magnetic reconnection as a rather robust, disc-internal ionisation channel that should be included in thermo-chemical and dynamical models of protoplanetary discs.
@@ 1 Introduction
Young, low-mass pre-main-sequence stars in the T Tauri phase, typically a few million years old, continue to contract and to accrete material from their circumstellar discs. How they remove the excess angular momentum to enable accretion remains an open problem (Venuti et al., 2014; Hartmann et al., 2016; Manara et al., 2016). Most viable scenarios invoke magnetic fields coupled to partly ionised gas, through either turbulent torques driven by (magneto)hydrodynamic (M)HD instabilities (Shakura and Sunyaev, 1973), or laminar torques associated with magnetised winds and jets (Blandford and Payne, 1982; Jacquemin-Ide et al., 2019). In this work, we focus on the consequences of MHD turbulence, where present in the disc. We restrict our attention to protoplanetary discs where the magneto-rotational instability (MRI) is likely the primary turbulence driver (Balbus and Hawley, 1991; Balbus, 2003). Crucially, in any accretion scenario where magnetic stresses regulate angular momentum transport, magnetic energy must be dissipated. In MRI-active layers, current sheets form so that dissipation through magnetic reconnection is a natural by-product of turbulent, magnetised discs (Pucci et al., 2024).
Reconnection can occur in several young stellar object (YSO) environments: in stellar coronae during flares (Benz, 2017), in the star–disc interaction region near the truncation radius (Ferreira et al., 2000; Fatuzzo et al., 2023), and, central to this paper, within the disc’s MRI-active layers, where field-line turbulence continually regenerates current sheets (de Gouveia Dal Pino et al., 2010). Besides, sustaining MRI requires a minimum ionisation level. In YSO environments several sources of ionisation have been already investigated: UV and X-ray radiation either coming directly from the star or from the background radiation field (Glassgold et al., 1997), Galactic Cosmic Ray (GCRs), (Cleeves et al., 2013b; Padovani et al., 2018), radionuclide ionisation (Cleeves et al., 2013a), stellar energetic particles (SEP ; Rab et al. 2017; Rodgers-Lee et al. 2017) and in-situ accelerated particles, in shocks (Padovani et al., 2016; Offner et al., 2019) or by magnetic reconnection in flares (Brunn et al., 2023, 2024). In the latter work, reconnection-accelerated EPs were shown to be an exceptionally efficient ionisation source, exceeding all others in their injection region (inner disc, $R\lesssim 1$ au), which motivated the present work.
@@
As MHD turbulence and magnetic reconnection develop in accretion discs, a part of the stored magnetic energy is possibly released in the form of non-thermal particles. Magnetic reconnection sites are known to accelerate non-thermal particles (see reviews in different astrophysical or space plasmas contexts by Marcowith et al. 2020; Oka et al. 2023; Guo et al. 2024). Particles can be accelerated in different variety of processes: Fermi-I like acceleration associated to the electric field induced in gas inflows (de Gouveia dal Pino and Lazarian, 2005; Drury, 2012), or due to curved magnetic fields. Acceleration can be due to the betatron effect in increasing magnetic zones under the conservation of the adiabatic particle moment, in magnetic contracting islands (Oka et al., 2010) or by direct electric field acceleration (Oka et al., 2023). Whatever the exact acceleration process, magnetic reconnection can be seen as a mechanism able to inject non-thermal particles into large scale turbulent flows as magnetic reconnection and turbulence are intimately inter-related.
@@
The paper is organised as follows. Sec. 2 outlines the disc thermo-chemical setup and links it to turbulence and reconnection scalings; Sec. 3 introduces our EP injection/acceleration prescription and transport setup; Sec. 4 presents ionisation (and associated heating) maps and identifies where EPs dominate over X-rays and GCRs; Sec. 5 examines implications and limitations; Sec. 6 summarises the main findings and outlines future directions.
2 Physical framework – From thermochemistry to turbulence-driven reconnection
@@
In this section we introduce the physical framework. We first outline the disc thermo-chemical background, then summarise observational and theoretical constraints on turbulence, and finally present the reconnection model that sets the parameters for EP acceleration. The left panel of Fig. 1 sketches our scenario: a protoplanetary disc with a weakly ionised, MRI-stable mid-plane overlain by a magnetically coupled atmosphere where sub-Alfvénic MRI turbulence bends field lines and builds current sheets. Magnetic reconnection in these atmospheric layers accelerates supra-thermal particles that propagate quasi-ballistically along near-vertical field lines  [^1] In this work we consider by default the case of vertical field lines transiting from the ionised layer toward the neutral disc, more complex magnetic geometries will be addressed in a forthcoming article, see e.g. Brunn et al. (2023) for a discussion of magnetic topologies in accretion discs. towards the mid-plane. This section presents the quantitative arguments and assumptions that justify this scenario and defines the quantities used in the following model.
@@
Figure 1: Left: schematic radial cut from $R\sim 0.1$ to $\sim 50$ AU presenting our scenario qualitatively. The disc is vertically stratified into a neutral mid-plane (blue; MRI-stable), where energetic particles (EPs) arriving from above travel quasi-ballistically (CSDA-like) along near-vertical paths, and an ionised atmosphere (yellow; low-$\beta_{p}$ MRI-active), where sub-Alfvénic turbulence ($M_{A}<1$) triggers turbulent magnetic reconnection that injects EPs, which then stream downward (black arrows). In the bottom-left part, a patchy variant of the ionised atmospheric layer is sketched, with a lower covering area and possible lateral growth (black arrows). In the ionised turbulent layer, EP transport could be treated as stochastic. The inner star–disc interaction zone treated in Brunn et al. (2023, 2024) is indicated at $R\lesssim 0.3$. Right: radial cut from $R\sim 0.07$ to $\sim 600$ AU displaying the disc chemical and turbulent structure used in this work based on the ProDiMo model. The top panel shows the total hydrogen number density $n_{\rm H,tot}$ (colour) with vertical column-density contours $N=10^{15},10^{19},10^{21},10^{24}\,\mathrm{cm^{-2}}$ (black dashed). The red dashed line marks the ionised–atomic transition ($\mathrm{H^{+}/H}$) defining the boarder between the region where $\mathrm{H^{+}}$ dominates and the region where $\mathrm{H}$ dominates. Analogously, the orange dashed line delimit the atomic–molecular transition ($\mathrm{H/H_{2}}$). The bottom panel displays the Alfvénic Mach number $M_{A}\equiv v_{\rm turb}/V_{A}$ (colour). The red curve traces $M_{A}=1$. The sub-Alfvénic region ($M_{A}<1$) identifies where fast, turbulence-enabled reconnection, thus EP acceleration is expected. The solid red lines delimit the region where EP are efficiently accelerated, i.e. to energy more than 10 MeV. The white region is the so called ”dead zone”, where MRI is inefficient (see criterion Eq. 23), there is no turbulence, $M_{A}$ is set to 0.
@@ 2.1 Protoplanetary Disc Model (ProDiMo )
ProDiMo [^2] https://prodimo.iwf.oeaw.ac.at revision: 66efbd75 (2023/06/27) is a 2D radiation thermo-chemical code for protoplanetary discs (Woitke et al., 2009; Kamp et al., 2010; Thi et al., 2011; Woitke et al., 2016). It self-consistently couples disc structure, wavelength-dependent radiative transfer, gas/dust chemistry, and thermal balance to predict physical conditions and synthetic spectra over a broad wavelength range. The disc is treated on an axisymmetric grid (R,Z) for $R\approx 0.07-600$ AU and Z/R=00.5Z/R=0-0.5 with a resolution set to $100\times 100$. Frequency-dependent dust continuum radiative transfer yields the local radiation field and dust temperature, including stellar and interstellar UV and stellar X-rays, enabling photochemistry and heating/cooling in the atmosphere. The chemical network spans 235 species and 3143 reactions (Kamp et al., 2017; Rab et al., 2017), with ice freeze-out and grain desorption (Thi et al., 2011). Heating processes include photoelectric emission, X-ray/Cosmic-Ray ionisation, viscous and collisional heating; cooling is treated via non-LTE line emission (e.g. CO, Fe II, O I), gas–dust coupling, and molecular lines. Iterative coupling ensures a consistent vertical structure (Arabhavi et al., 2022).
@@
We adopt the fiducial T Tauri disc of Rab et al. (2017), with $M_{\star}=0.7M_{\odot}$, $T_{\star}=4000\mathrm{K}$, $L_{\star}=1.0L_{\odot}$, and an accretion rate $\dot{M}=10^{-8}M_{\odot}\mathrm{yr}^{-1}$, as detailed in their Table 4, for our thermo-chemical background. This setup has been widely used (Woitke et al., 2016; Kamp et al., 2017; Woitke et al., 2019). From the model, we extract at each grid cell $(R,Z)$: the total hydrogen nuclei density, $n_{\rm H,tot}=n_{\rm H}+2n_{\rm H_{2}}$; the gas temperature $T$; the electron fraction $x_{e}$; the local composition ($n_{\rm H}$, $n_{\rm H_{2}}$, $n_{\rm He}$) used to characterize the turbulence and to build mixture-averaged energy-loss functions for electrons and protons; and the vertical column densities $N(R,Z)\equiv\int_{0.5R}^{Z}n_{\rm H,tot}(R,z)dz$ required for the transport model. These quantities set the injection EP energetics (Sect.3), determine energy losses during propagation (Sect.4), and yield the reference X-ray and CR ionisation rates used for comparison.
@@
The top-right panel of Fig. 1 display the disc structure used in this work, as per ProDiMo , it shows a colour map of the total hydrogen number density, $n_{\mathrm{H,tot}}$. Black dashed contours trace the vertical column density $N$ at $10^{15},\,10^{19},\,10^{21}$, and $10^{24}\,\mathrm{cm^{-2}}$. The red dashed curve marks the $\mathrm{H^{+}/H}$ transition (ionised–atomic boundary), and the orange dashed curve marks the $\mathrm{H/H_{2}}$ transition (atomic–molecular boundary). In the following, we adopt the following notation to designate the different regions of the disc: (i) Atmosphere: $10^{15}\leq N<10^{19}\,\mathrm{cm^{-2}}$, highly ionised. (ii) Surface: $10^{19}\lesssim N\lesssim 10^{21}\,\mathrm{cm^{-2}}$, predominantly atomic gas. (iii) Intermediate layer: $10^{21}\lesssim N\lesssim 10^{24}\,\mathrm{cm^{-2}}$, dominated by molecular gas. (iv) Deep layer: $N\gtrsim 10^{24}\,\mathrm{cm^{-2}}$, characterised by strong visual extinction ($A_{V}\gtrsim 10)$.
@@ 2.2.1 Observational constraints on turbulence in discs
Quantifying turbulence in discs relies primarily on (i) spectrally resolved molecular line observations that separate thermal from non-thermal broadening, (ii) the use of multiple tracers and isotopologues that probe different heights, and (iii) indirect constraints from dust settling. High spectral–spatial resolution ALMA data of CO and its isotopologues generally indicate weak turbulence in many systems, particularly in the outer disc ($R\gtrsim 30$ AU) and near the mid-plane, with typical non-thermal velocities bounded by $varv_{\rm turb}\lesssim 0.1c_{s}$ at radii of a few tens of AU (e.g. Flaherty et al., 2017; Rosotti, 2023). Independent inferences from the thickness of dust layers likewise point to low turbulent stirring. For instance, in HL Tau, dust vertical stratification implies $varv_{\rm turb}\sim 10^{-2}c_{s}$ at $R\sim 100$ AU (Pinte et al., 2016). However, a contrasting picture emerges in some sources and at higher altitudes above the mid-plane. Broader CO lines in DM Tau suggest $varv_{\rm turb}\sim 0.2\text{--}0.3c_{s}$ at $R\sim 20\text{--}50$ AU (Flaherty et al., 2020). Moreover, chemically selective tracers that originate in the upper layers (e.g. CN, C2H) reveal non-negligible turbulence at $Z/R\gtrsim 0.2\text{--}0.3$ and large radii in IM Lup (Paneque-Carreño et al., 2024). Recent ALMA CO analyses also in IM Lup, report on $varv_{\rm turb}\sim 0.18\text{--}0.30c_{s}$ at $R\sim 60$ AU (Flaherty et al., 2024). Taken together, these measurements suggest a vertical gradient in the turbulent amplitude: turbulence is generally weakest near the mid-plane and stronger in the disc atmosphere, consistent with expectations for more ionised, MRI-active surface layers. This work focuses on a radial domain in the inner disc ($R\simeq 0.3$ to $30$ AU) that is not constrained by ALMA. The turbulence is unknown. However, it is expected that within this region MRI is active (Lesur et al., 2023). The inner boundary reflects the zone of direct stellar magnetospheric influence due to stellar flaring loops, typically extending to $R\sim 0.3-0.5$ AU (Brunn et al., 2024).

### v2

# Energetic particles accelerated via turbulent magnetic reconnection in protoplanetary discs – I. Ionisation rates

@@ Abstract
Context. Ionisation controls the chemistry, thermal balance, and magnetic coupling in protoplanetary discs. However, standard ionisation vectors such as stellar UV, X-rays, Galactic cosmic rays might not be efficient enough, as UV/X-rays are attenuated rapidly with depth, while Galactic cosmic rays are modulated. Turbulence-induced magnetic reconnection in disc atmospheric layers offers a physically motivated, in situ source of energetic particles (EPs) that has never been considered.
@@
Methods. We adopt a Fermi-like acceleration model with parameters linked to a turbulent reconnection geometry trigger by the magneto-rotational instability, yielding a steady-state energy distribution of the EP forming a power-law of index $p=2.5$. We propagate electrons and protons through the disc and compute primary and secondary ionisation and associated heating on a fiducial T Tauri disc model background. The non-thermal normalisation is set by the fraction of local viscous accretion energy dissipation channelled to EPs, parametrised by $\kappa$.
Results. For $\kappa\gtrsim 0.4\%$, EPs ionisation overpass standard sources such as X-rays and Galactic cosmic rays in the disc atmosphere and intermediate/deep layers out to radii of a few tens of astronomical units. Even at $\kappa\sim 0.025\%$, EPs contribute at the few-percent level, thus are chemically and dynamically relevant. The EP-induced heating complements UV/X-ray heating in the atmosphere and persists deeper. These results identify EPs accelerated by turbulence-induced magnetic reconnection as a rather robust, disc-internal ionisation channel that should be included in thermo-chemical and dynamical models of protoplanetary discs.
@@ 1 Introduction
Young, low-mass pre-main-sequence stars in the T Tauri phase, typically a few million years old, continue to contract and to accrete material from their circumstellar discs. How they remove the excess angular momentum to enable accretion remains an open problem (Venuti et al., 2014; Hartmann et al., 2016; Manara et al., 2016). Most viable scenarios invoke magnetic fields coupled to partly ionised gas, through either turbulent torques driven by magneto-hydrodynamic (MHD) instabilities (Shakura and Sunyaev, 1973), or laminar torques associated with magnetised winds and jets (Blandford and Payne, 1982; Jacquemin-Ide et al., 2019). In this work, we focus on the consequences of MHD turbulence, where present in the disc. We restrict our attention to protoplanetary discs where the magneto-rotational instability (MRI) is likely the primary turbulence driver (Balbus and Hawley, 1991; Balbus, 2003). Crucially, in any accretion scenario where magnetic stresses regulate angular momentum transport, magnetic energy must be dissipated. In MRI-active layers, current sheets form so that dissipation through magnetic reconnection is a natural by-product of turbulent, magnetised discs (Pucci et al., 2024).
Reconnection can occur in several young stellar object (YSO) environments: in stellar coronae during flares (Benz, 2017), in the star–disc interaction region near the truncation radius (Ferreira et al., 2000; Fatuzzo et al., 2023), and, central to this paper, within the disc’s MRI-active layers, where field-line turbulence continually regenerates current sheets (de Gouveia Dal Pino et al., 2010). Besides, sustaining MRI requires a minimum ionisation level. In YSO environments several sources of ionisation have already been investigated: UV and X-ray radiation either coming directly from the star or from the background radiation field (Glassgold et al., 1997), Galactic cosmic rays (GCRs), (Cleeves et al., 2013b; Padovani et al., 2018), radionuclide ionisation (Cleeves et al., 2013a), stellar energetic particles (SEP ; Rab et al. 2017; Rodgers-Lee et al. 2017) and in-situ accelerated particles, in shocks (Padovani et al., 2016; Offner et al., 2019) or by magnetic reconnection in flares (Brunn et al., 2023, 2024). In the latter work, reconnection-accelerated EPs were shown to be an exceptionally efficient ionisation source, exceeding all others in their injection region (inner disc, $R\lesssim 1$ au), which motivated the present work.
@@
As MHD turbulence and magnetic reconnection develop in accretion discs, a part of the stored magnetic energy is possibly released in the form of non-thermal particles. Magnetic reconnection sites are known to accelerate non-thermal particles (see reviews in different astrophysical or space plasmas contexts by Marcowith et al. 2020; Oka et al. 2023; Guo et al. 2024). Particles can be accelerated in a variety of processes: Fermi-I like acceleration associated with the electric field induced in gas inflows (de Gouveia dal Pino and Lazarian, 2005; Drury, 2012), or due to curved magnetic fields. Acceleration can be due to the betatron effect in increasing magnetic zones under the conservation of the adiabatic particle moment, in magnetic contracting islands (Oka et al., 2010) or by direct electric field acceleration (Oka et al., 2023). Whatever the exact acceleration process, magnetic reconnection can be seen as a mechanism able to inject non-thermal particles into large scale turbulent flows as magnetic reconnection and turbulence are intimately inter-related.
@@
The paper is organised as follows. Sec. 2 outlines the disc thermo-chemical set-up and links it to turbulence and reconnection scalings. Sec. 3 introduces our EP injection and acceleration prescription as well as the transport set-up. Sec. 4 presents ionisation (and associated heating) maps and identifies where EPs dominate over X-rays and GCRs. Sec. 5 examines implications and limitations. Sec. 6 summarises the main findings and outlines future directions.
2 Physical framework: From thermochemistry to turbulence-driven reconnection
@@
In this section we introduce the physical framework. We first outline the disc thermo-chemical background, then summarise observational and theoretical constraints on turbulence, and finally present the reconnection model that sets the parameters for EP acceleration. The left panel of Fig. 1 sketches our scenario: a protoplanetary disc with a weakly ionised, MRI-stable mid-plane overlain by a magnetically coupled atmosphere where sub-Alfvénic MRI turbulence bends field lines and builds current sheets. Magnetic reconnection in these atmospheric layers accelerates supra-thermal particles that propagate quasi-ballistically along near-vertical field lines  [^1] In this work we consider by default the case of vertical field lines transiting from the ionised layer towards the neutral disc. More complex magnetic geometries will be addressed in a forthcoming article (see e.g. Brunn et al. 2023 for a discussion of magnetic topologies in accretion discs). towards the mid-plane. This section presents the quantitative arguments and assumptions that justify this scenario and defines the quantities used in the following model.
@@
Figure 1: Left: Schematic radial cut from $R\sim 0.1$ to $\sim 50$ AU presenting our scenario qualitatively. The disc is vertically stratified into a neutral mid-plane (blue; MRI-stable), where EPs arriving from above travel quasi-ballistically (CSDA-like) along near-vertical paths, and an ionised atmosphere (yellow; low-$\beta_{p}$ MRI-active), where sub-Alfvénic turbulence ($M_{A}<1$) triggers turbulent magnetic reconnection that injects EPs, which then stream downwards (black arrows). In the bottom left part, a patchy variant of the ionised atmospheric layer is sketched, with a lower covering area and possible lateral growth (black arrows). In the ionised turbulent layer, EP transport could be treated as stochastic. The inner star–disc interaction zone treated in Brunn et al. (2023, 2024) is indicated at $R\lesssim 0.3$. Right: Radial cut from $R\sim 0.07$ to $\sim 600$ AU displaying the disc chemical and turbulent structure used in this work based on the ProDiMo model. The top panel shows the total hydrogen number density, $n_{\rm H,tot}$ (colour), with vertical column-density contours $N=10^{15},10^{19},10^{21},10^{24}\,\mathrm{cm^{-2}}$ (dashed black). The dashed red line marks the ionised–atomic transition ($\mathrm{H^{+}/H}$) defining the boarder between the region where $\mathrm{H^{+}}$ dominates and the region where $\mathrm{H}$ dominates. Analogously, the dashed orange line delimits the atomic–molecular transition ($\mathrm{H/H_{2}}$). The bottom panel displays the Alfvénic Mach number $M_{A}\equiv v_{\rm turb}/V_{A}$ (colour). The red curve traces $M_{A}=1$. The sub-Alfvénic region ($M_{A}<1$) identifies where fast, turbulence-enabled reconnection, and thus EP acceleration, is expected. The solid red lines delimit the region where EP are efficiently accelerated, i.e. to energy more than 10 MeV. The white region is the so-called ’dead zone’, where MRI is inefficient (see criterion Eq. 23), there is no turbulence, $M_{A}$ is set to 0.
@@ 2.1 Protoplanetary disc model (ProDiMo )
ProDiMo [^2] https://prodimo.iwf.oeaw.ac.at revision: 66efbd75 (2023/06/27) is a 2D radiation thermo-chemical code for protoplanetary discs (Woitke et al., 2009; Kamp et al., 2010; Thi et al., 2011; Woitke et al., 2016). It self-consistently couples disc structure, wavelength-dependent radiative transfer, gas and dust chemistry, and thermal balance to predict physical conditions and synthetic spectra over a broad wavelength range. The disc is treated on an axisymmetric grid (R,Z) for $R\approx 0.07-600$ AU and Z/R=00.5Z/R=0-0.5 with a resolution set to $100\times 100$. Frequency-dependent dust continuum radiative transfer yields the local radiation field and dust temperature, including stellar and interstellar UV and stellar X-rays, enabling photochemistry and heating and cooling in the atmosphere. The chemical network spans 235 species and 3143 reactions (Kamp et al., 2017; Rab et al., 2017), with ice freeze-out and grain desorption (Thi et al., 2011). Heating processes include photoelectric emission, X-ray/cosmic-ray ionisation, viscous and collisional heating; cooling is treated via non-local thermal equilibrium line emission (e.g. CO, Fe II, O I), gas–dust coupling, and molecular lines. Iterative coupling ensures a consistent vertical structure (Arabhavi et al., 2022).
@@
We adopted the fiducial T Tauri disc of Rab et al. (2017), with $M_{\star}=0.7M_{\odot}$, $T_{\star}=4000\mathrm{K}$, $L_{\star}=1.0L_{\odot}$, and an accretion rate of $\dot{M}=10^{-8}M_{\odot}\mathrm{yr}^{-1}$, as is detailed in their Table 4, for our thermo-chemical background. This set-up has been widely used (Woitke et al., 2016; Kamp et al., 2017; Woitke et al., 2019). From the model, we extracted at each grid cell $(R,Z)$: the total hydrogen nuclei density, $n_{\rm H,tot}=n_{\rm H}+2n_{\rm H_{2}}$; the gas temperature, $T$; the electron fraction, $x_{e}$; the local composition ($n_{\rm H}$, $n_{\rm H_{2}}$, $n_{\rm He}$) used to characterise the turbulence and to build mixture-averaged energy-loss functions for electrons and protons; and the vertical column densities, $N(R,Z)\equiv\int_{0.5R}^{Z}n_{\rm H,tot}(R,z)dz$, required for the transport model. These quantities set the injection EP energetics (Sect.3), determine energy losses during propagation (Sect.4), and yield the reference X-ray and CR ionisation rates used for comparison.
@@
The top right panel of Fig. 1 displays the disc structure used in this work. As per ProDiMo , it shows a colour map of the total hydrogen number density, $n_{\mathrm{H,tot}}$. Dashed black contours trace the vertical column density, $N$, at $10^{15},\,10^{19},\,10^{21}$, and $10^{24}\,\mathrm{cm^{-2}}$. The dashed red curve marks the $\mathrm{H^{+}/H}$ transition (ionised–atomic boundary), and the dashed orange curve marks the $\mathrm{H/H_{2}}$ transition (atomic–molecular boundary). In the following, we adopt the following notation to designate the different regions of the disc: (i) Atmosphere: $10^{15}\leq N<10^{19}\,\mathrm{cm^{-2}}$, highly ionised. (ii) Surface: $10^{19}\lesssim N\lesssim 10^{21}\,\mathrm{cm^{-2}}$, predominantly atomic gas. (iii) Intermediate layer: $10^{21}\lesssim N\lesssim 10^{24}\,\mathrm{cm^{-2}}$, dominated by molecular gas. (iv) Deep layer: $N\gtrsim 10^{24}\,\mathrm{cm^{-2}}$, characterised by strong visual extinction ($A_{V}\gtrsim 10)$.
@@ 2.2.1 Observational constraints on turbulence in discs
Quantifying turbulence in discs relies primarily on (i) spectrally resolved molecular line observations that separate thermal from non-thermal broadening, (ii) the use of multiple tracers and isotopologues that probe different heights, and (iii) indirect constraints from dust settling. High spectral–spatial resolution ALMA data of CO and its isotopologues generally indicate weak turbulence in many systems, particularly in the outer disc ($R\gtrsim 30$ AU) and near the mid-plane, with typical non-thermal velocities bounded by $varv_{\rm turb}\lesssim 0.1c_{s}$ at radii of a few tens of astronomical units (e.g. Flaherty et al., 2017; Rosotti, 2023). Independent inferences from the thickness of dust layers likewise point to low turbulent stirring. For instance, in HL Tau, dust vertical stratification implies $varv_{\rm turb}\sim 10^{-2}c_{s}$ at $R\sim 100$ AU (Pinte et al., 2016). However, a contrasting picture emerges in some sources and at higher altitudes above the mid-plane. Broader CO lines in DM Tau suggest $varv_{\rm turb}\sim 0.2\text{--}0.3c_{s}$ at $R\sim 20\text{--}50$ AU (Flaherty et al., 2020). Moreover, chemically selective tracers that originate in the upper layers (e.g. CN, C2H) reveal non-negligible turbulence at $Z/R\gtrsim 0.2\text{--}0.3$ and large radii in IM Lup (Paneque-Carreño et al., 2024). Recent ALMA CO analyses also in IM Lup, report on $varv_{\rm turb}\sim 0.18\text{--}0.30c_{s}$ at $R\sim 60$ AU (Flaherty et al., 2024). Taken together, these measurements suggest a vertical gradient in the turbulent amplitude: turbulence is generally weakest near the mid-plane and stronger in the disc atmosphere, consistent with expectations for more ionised, MRI-active surface layers. This work focuses on a radial domain in the inner disc ($R\simeq 0.3$ to $30$ AU) that is not constrained by ALMA. The turbulence is unknown. However, it is expected that within this region MRI is active (Lesur et al., 2023). The inner boundary reflects the zone of direct stellar magnetospheric influence due to stellar flaring loops, typically extending to $R\sim 0.3-0.5$ AU (Brunn et al., 2024).